### v1

# What Makes a Good Doctor Response? An Analysis on a Romanian Telemedicine Platform

Abstract
Text-based telemedicine has become a common mode of care, requiring clinicians to deliver medical advice clearly and effectively in writing. As platforms increasingly rely on patient ratings and feedback, clinicians face growing pressure to maintain satisfaction scores, even though these evaluations often reflect communication quality more than clinical accuracy. We analyse patient satisfaction signals in Romanian text-based telemedicine. Using a sample of 77,334 anonymised patient question–doctor response pairs, we model feedback as a binary outcome, treating thumbs-up responses as positive and grouping negative or absent feedback into the other class. We extract interpretable, predominantly language-agnostic features (e.g., length, structural characteristics, readability proxies), along with Romanian LIWC psycholinguistic features and politeness/hedging markers where available. We train a classifier with a time-based split and perform SHAP-based analyses, which indicate that patient and clinician history features dominate prediction, functioning as strong priors, while characteristics of the response text provide a smaller but, crucially, actionable signal. In subgroup correlation analyses, politeness and hedging are consistently positively associated with patient feedback, whereas lexical diversity shows a negative association.
Keywords: telemedicine; patient feedback; doctor–patient communication; Romanian NLP; interpretability; SHAP; psycholinguistic features;
[…]
What Makes a Good Doctor Response?
An Analysis on a Romanian Telemedicine Platform
Adrian Cosma, Cosmin Dumitrache, Emilian Radoi
IDSIA, Dalle Molle Institute for Artificial Intelligence, Lugano, CH
University Politehnica of Bucharest, Bucharest, RO
Correspondence: emilian.radoi@upb.ro
Abstract content
1.  Introduction
Text-based telemedicine has quickly evolved from a niche offering to a common way people receive care. As more consultations move online, clinicians are increasingly practising medicine in writing. They must give clinically sound advice while explaining it clearly, using language that patients can easily understand and act upon. This change also introduces expectations that have emerged from other online services, especially regarding transparency and feedback. Patients are accustomed to ratings and reviews in other digital services, and many now expect similar indicators when deciding whether to trust a clinician on a telemedicine platform.
As a result, many telemedicine platforms now collect structured feedback after each consultation (e.g., thumbs up / thumbs down or star ratings, optionally accompanied by free-text comments). In contrast to in-person care, where feedback is relatively rare and often informal, online platforms can generate large volumes of explicit evaluations for individual clinicians.
While review systems have become ubiquitous across online platforms, applying them to medical consultations presents unique challenges. Doctors find themselves evaluated not by medical peers, but by patients who generally lack the clinical expertise to assess medical accuracy. This creates a fundamental tension: patient satisfaction and review scores are often driven more by how medical advice is communicated and framed than by its clinical correctness. Consequently, doctors face pressure to maintain high ratings in order to sustain patient demand, yet they lack clear guidance on which aspects of their communication to change. This creates both an opportunity and a challenge: feedback can incentivize better patient communication and help platforms improve quality, but it can also be difficult for doctors to interpret and act upon, especially when it is negative.
This work attempts to address a key question in telemedicine platforms: which parts of a doctor’s written response are most associated with higher patient satisfaction, and do these patterns differ across patient and clinician subgroups?
A further motivation for this study is the linguistic context. We analyse interactions written in Romanian Niculae et al. (2025b); Rogoz et al. (2025), a low-resource languageNigatu et al. (2024). Most NLP research in healthcare has focused on English and other high-resource languages, with relatively limited work on Romanian. As a result, key NLP resources for Romanian, such as clinical-domain models and standardised preprocessing pipelines, remain less developed and less thoroughly validated Tufiș and Barbu Mititelu (2014); Dumitrescu et al. (2018). To keep the analysis interpretable and realistic, we therefore rely primarily on language-agnostic, surface-level measurements (e.g., length and structure) and incorporate Romanian-specific resources through the Romanian LIWC lexicon Dudău and Sava (2022).
In this work, we use fully anonymised real-world data consisting of Romanian-language text-based medical consultations, to study which factors correlate with patient satisfaction, as reflected in post-consultation ratings. We treat feedback as the result of multiple interacting components:
  1. (i)
    Patient profile and prior behaviour on the platform
  2. (ii)
    Clinician profile and platform activity
  3. (iii)
    Question metadata (e.g., time, medical specialization, inter alia)
  4. (iv)
    Properties of the written response
Importantly, we focus on observable, surface-level attributes of doctors’ responses, features that can be measured and acted upon, rather than attempting to assess medical correctness.
We extract features related to readability and form (e.g., length and structure), psycholinguistic signals (e.g., LIWC categories), and metadata linked to the patient, doctor and question. We then train a gradient-boosting model to predict patient feedback, choosing it as a strong and simple baseline that supports post-hoc interpretability. To understand what the model learns, we use SHAP Lundberg and Lee (2017) to quantify global feature importance and provide local explanations for individual interactions, highlighting which response characteristics most strongly drive predicted satisfaction and how these effects vary across patient and doctor subgroups.
2.  Related Work
Large language models (LLMs) are increasingly adapted to medical settings via pre-training, fine-tuning, and prompting on clinical corpora Zhou et al. (2023). Much of the literature evaluates exam-style or multiple-choice QA, where strong performance can be achieved under appropriate adaptation strategies Zhou et al. (2023). More recent "copilot" frameworks aim to support broader clinical workflows without task-specific fine-tuning Ren et al. (2024), but open-ended, real-world interaction settings remain less studied.
Patient-facing deployment depends not only on correctness but also on perceived reliability and empathy. AI-labeled medical advice can be rated as less trustworthy or empathetic, lowering willingness to follow it even when content is held constant Reis et al. (2024); Hohenstein et al. (2023). Conversely, users judge AI responses as comparable to clinicians despite low accuracy Shekar et al. (2024). Reliability can also be user-dependent: model behavior may degrade in public-facing use, plausibly because outputs adapt to the quality and specificity of user inputs Team ; Bean et al. (2025).
Beyond LLMs, medical humanities and sociolinguistics emphasize that patients value empathy, clarity, attention to lived experience, and appropriate boundary-setting Stergiopoulos and Martimianakis (2023). A well presented medical advice is more likely to have a positive impact Martin et al. (2005) and to attract positive reviews, directly impacting doctor reputation Street Jr et al. (2009); Sargeant et al. (2008). Quantitative work links physicians’ language to outcomes in online consultations, for example positive-emotion and instrumental language correlating with higher returns Geng et al. (2024). Classic analyses also report systematic gender differences in doctor–patient dialogue Meeuwesen et al. (1991), motivating subgroup analyses when studying satisfaction.
Rather than asking whether LLMs provide correct answers, we study what patients reward in human clinician writing. Using a large Romanian telemedicine dataset (77k+ question–answer pairs), we measure which observable properties of replies (surface form, readability proxies, psycholinguistic features) are associated with positive feedback. This complements Romanian-focused efforts on optimizing LLM-assisted doctor communication that rely on smaller exploratory datasets Niculae et al. (2025a). Our goal is not to assess clinical correctness, but to provide an interpretable view of communication cues that correlate with satisfaction.
3.  Method
We study text-based consultations consisting of patient questions paired with doctor responses (N=77,334N{=}77{,}334). After an interaction, patients may leave structured feedback (thumbs-up / thumbs-down), but in practice most interactions receive no explicit rating and explicit negative feedback is rare. We therefore model patient satisfaction as a binary prediction task, treating thumbs-up as the positive class and grouping thumbs-down and missing feedback into the other class. The resulting dataset is approximately balanced (39,201 positive vs. 38,133 other). Although fully anonymised, the data is private and cannot be released.
A particular aspect of the studied text consultations is that doctors may address clarification questions before providing an answer. We retain this structure and distinguish between (i) Simple Questions and (ii) Questions with Clarifications (a clarification step occurs), which we also use for stratified reporting.
Feature extraction To keep the analysis interpretable and robust in a lower-resource language setting, we focus primarily on surface-level, mostly language-agnostic features, and add Romanian-specific resources where available. We extract features from four sources: (i) patient profile – platform tenure, activity, and prior behaviour (e.g., historical thumbs-up rate, number of posed questions per year), (ii) doctor profile – metadata (e.g., workplace, education) and platform activity and historical feedback rate. (iii) question metadata and patient text – basic structure and readability proxies (e.g., length), plus LIWC psycholinguistic categories, and (iv) doctor response text – length / structure (characters, tokens, punctuation counts such as question marks), concreteness proxies (e.g., number of digits), and lexicon-based markers (e.g., politeness, hedging, recommendation cues), as well as LIWC features. We compute a fixed feature vector of 250 features spanning metadata, history and text-derived signals.
Prediction model and evaluation protocol We train a gradient-boosted decision tree classifier for binary feedback prediction. To reduce temporal leakage, we use a time-based split: training uses earlier interactions and evaluation is performed on later interactions, for a 80:20 train:test split. We report performance both on the full validation set and on relevant subsets.
Model interpretation and subgroup analyses To understand which factors drive predictions, we compute SHAP values on the trained model. We summarize global importance with mean absolute SHAP values and use local explanations to inspect individual decisions. In addition, to probe whether associations differ across populations, we perform subgroup correlation analyses between selected text features and feedback, stratifying by patient gender (patient preferences) and doctor gender (doctor behavior), and report the resulting patterns descriptively.
4.  Results
Subset Split ROC-AUC PR-AUC F1 Precision Recall
Simple Questions Val 0.6780 0.6865 0.6206 0.6206 0.6209
Only Clarifications Val 0.7153 0.5857 0.6367 0.6411 0.6493
All Questions Val 0.7192 0.6313 0.6518 0.6542 0.6577
New Patients Val 0.6479 0.6003 0.5804 0.6061 0.5963
Old Patients Val 0.7656 0.6533 0.6951 0.6968 0.6937
Simple Questions Val High-conf 0.6951 0.6138 0.6456 0.7724 0.6416
Only Clarifications Val High-conf 0.8348 0.6542 0.7751 0.7960 0.7603
All Questions Val High-conf 0.8357 0.6987 0.7889 0.8001 0.7806
Table 1: Feedback prediction performance across question subsets and patient tenure groups (New vs. Old), reported on the validation split and a high-confidence subset.
Figure 1: Mean absolute SHAP values for the CatBoost feedback classifier trained on 250 features spanning patient profile/history, doctor profile/history and text-derived features from both sides. Larger values indicate greater impact on predictions (magnitude only; SHAP sign is not shown).
Figure 2: Pearson correlations between selected doctor-response text features and positive feedback, aggregated over all interactions. Positive values indicate features that tend to co-occur with thumbs-up feedback, while negative values indicate features that tend to co-occur with negative or missing feedback.
In Table 1, we show classifier performance across subgroups. Since the classifier is properly calibrated (calibration error of 1.5%), high confidence predictions have high performance in terms of ROC-AUC score. Results are also better for clarification-type questions: clarification questions may function as a visible signal of attention and engagement.
Correlation analysis by doctor and patient gender. Figure 2 shows Pearson correlations between doctor-text features and binary positive feedback, computed separately by doctor gender (Doctor Behaviour) and by patient gender (Patient Preferences). Across all splits, the largest positive correlations are for politeness markers and hedging, consistent with prior work Stergiopoulos and Martimianakis (2023). Smaller positive correlations appear for features that often signal concrete, actionable replies, such as the number of digits, explicit recommendations, and question marks. The most consistent negative correlation is for lexical diversity, with additional negative correlations for several LIWC categories (e.g., liwc_work, liwc_affiliation, liwc_social). Higher lexical diversity can reflect more complex, less standardized phrasing, which may increase cognitive load; conversely, simpler, more formulaic language may read as clearer and safer Stergiopoulos and Martimianakis (2023).
The overall pattern is similar across groups, suggesting that both male and female patients tend to respond positively to polite, cautious, and concrete messages. Differences by patient gender are small. Differences by doctor gender are more visible: for several of the top positive features (especially politeness, hedging and digits), correlations are larger within the subset of replies written by female doctors than within those written by male doctors. Correlation magnitudes are modest overall (roughly $|r|\approx 0.02$$0.15$), so these results should be treated as descriptive.
Model explanations using SHAP. Figure 1 reports mean absolute SHAP values (feature importance). Overall, the model relies most on patient history features. The largest contributions come from the previous thumbs-up rate and the previous number of questions. These features mainly act as strong priors: patients with a higher past tendency to give positive feedback (or higher platform activity/tenure) are easier to predict.
Doctor-level features contribute less than patient-level history. The top doctor predictors are previous received thumbs-up rate, doctor workplace, number of previously helped patients, and number of years on the platform, along with past positive / negative feedback rates. This suggests the model captures persistent differences between clinicians and settings (e.g., workplace) beyond what is in the text.
Text features have smaller mean $|\mathrm{SHAP}|$ values than profile/history features, indicating that metadata dominates the prediction. Within text, doctor-side features are more influential than patient-side features. The most important doctor-text signals include number of questions, number of digits present, and several LIWC categories (e.g., liwc_work, liwc_negemo, liwc_time), as well as length proxies (n_characters, n_tokens). These patterns are consistent with feedback being associated with replies that are more interactive (questions/clarifications), more concrete (numbers) and more detailed (longer responses), confirming prior work Geng et al. (2024). Digits often encode dosage, timing, thresholds or step-by-step plans, which are surface markers of actionability. Patient-text features have comparatively small contributions. Satisfaction signals are identity- and expectation-loaded: many ratings reflect the rater’s baseline tendency and the clinician’s reputation/context more than the linguistic content. This is consistent with "perception-first" evaluation dynamics discussed in prior work Shekar et al. (2024), and motivates controlling for priors before making communication recommendations.
5.  Conclusions
We analysed 77k anonymised Romanian patient question–doctor response pairs to examine which characteristics of written clinician replies are associated with positive patient feedback in text-based telemedicine consultations. The main finding is that satisfaction is driven largely by stable priors (patient and doctor history), so predicting feedback mostly captures who tends to rate positively and who tends to receive positive ratings. Still, the response text contains smaller but actionable signals: politeness and calibrated uncertainty are consistently associated with positive feedback, as are concrete, action-oriented cues (e.g., digits and recommendations), while higher lexical diversity correlates negatively. Presumably, more high level and semantic response features metrics tailored to this specific use-case can be discovered automatically using feature discovery methods Cosma et al. (2026). These results suggest that patient feedback should be interpreted primarily as an incentive for clear, empathetic written communication, rather than as a proxy for medical accuracy.
  • A. Bean, R. E. Payne, G. Parsons, H. R. Kirk, J. Ciro, R. Mosquera-Gómez, S. Hincape, A. Ekanayaka, L. Tarassenko, L. Rocher, et al. (2025) Reliability of llms as medical assistants for the general public: a randomized preregistered study. Nature Medicine. Cited by: §2.
  • A. Cosma, O. Szehr, D. Kletz, A. Antonucci, and O. Pelletier (2026) Automatic prompt optimization for dataset-level feature discovery. External Links: 2601.13922, Link Cited by: §5.
  • D. P. Dudău and F. A. Sava (2022) The development and validation of the romanian version of linguistic inquiry and word count 2015 (ro-liwc2015). Current Psychology 41 (6), pp. 3597–3614. Cited by: §1.
  • S. D. Dumitrescu, A. M. Avram, L. Morogan, and S. Toma (2018) RoWordNet–a python api for the romanian wordnet. In 2018 10th International Conference on Electronics, Computers and Artificial Intelligence (ECAI), pp. 1–6. Cited by: §1.
  • S. Geng, Y. He, L. Duan, C. Yang, X. Wu, G. Liang, and B. Niu (2024) The association between linguistic characteristics of physicians’ communication and their economic returns: mixed method study. Journal of Medical Internet Research 26, pp. e42850. Cited by: §2, §4.
  • J. Hohenstein, R. F. Kizilcec, D. DiFranzo, Z. Aghajari, H. Mieczkowski, K. Levy, M. Naaman, J. Hancock, and M. F. Jung (2023) Artificial intelligence in communication impacts language and social relationships. Scientific reports 13 (1), pp. 5487. Cited by: §2.
  • S. M. Lundberg and S. Lee (2017) A unified approach to interpreting model predictions. CoRR abs/1705.07874. External Links: Link, 1705.07874 Cited by: §1.
  • L. R. Martin, S. L. Williams, K. B. Haskard, and M. R. DiMatteo (2005) The challenge of patient adherence. Therapeutics and clinical risk management 1 (3), pp. 189–199. Cited by: §2.
  • L. Meeuwesen, C. Schaap, and C. Van Der Staak (1991) Verbal analysis of doctor-patient communication. Social science & medicine 32 (10), pp. 1143–1150. Cited by: §2.
  • A. Niculae, A. Cosma, C. Dumitrache, and E. Radoi (2025a) Dr. copilot: a multi-agent prompt optimized assistant for improving patient-doctor communication in Romanian. In Proceedings of the 2025 Conference on Empirical Methods in Natural Language Processing: Industry Track, S. Potdar, L. Rojas-Barahona, and S. Montella (Eds.), Suzhou (China), pp. 1780–1792. External Links: Link, Document, ISBN 979-8-89176-333-3 Cited by: §2.
  • A. Niculae, A. Cosma, and E. Radoi (2025b) A retrieval-based approach to medical procedure matching in romanian. arXiv preprint arXiv:2503.20556. Cited by: §1.
  • H. H. Nigatu, A. L. Tonja, B. Rosman, T. Solorio, and M. Choudhury (2024) The Zeno’s Paradox of ‘Low-Resource’ Languages. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, Y. Al-Onaizan, M. Bansal, and Y. Chen (Eds.), Miami, Florida, USA, pp. 17753–17774. External Links: Document, Link Cited by: §1.
  • M. Reis, F. Reis, and W. Kunde (2024) Influence of believed ai involvement on the perception of digital medical advice. Nature Medicine 30 (11), pp. 3098–3100. Cited by: §2.
  • Z. Ren, Y. Zhan, B. Yu, L. Ding, and D. Tao (2024) Healthcare copilot: eliciting the power of general llms for medical consultation. arXiv preprint arXiv:2402.13408. Cited by: §2.
  • A. Rogoz, R. T. Ionescu, A. Anghel, I. Antone-Iordache, S. Coniac, and A. I. Ionescu (2025) MedQARo: a large-scale benchmark for evaluating large language models on medical question answering in romanian. arXiv preprint arXiv:2508.16390. Cited by: §1.
  • J. Sargeant, K. Mann, D. Sinclair, C. Van der Vleuten, and J. Metsemakers (2008) Understanding the influence of emotions and reflection upon multi-source feedback acceptance and use. Advances in Health Sciences Education 13, pp. 275–288. Cited by: §2.
  • S. Shekar, P. Pataranutaporn, C. Sarabu, G. A. Cecchi, and P. Maes (2024) People over trust ai-generated medical responses and view them to be as valid as doctors, despite low accuracy. arXiv preprint arXiv:2408.15266. Cited by: §2, §4.
  • E. Stergiopoulos and M. A. T. Martimianakis (2023) What makes a ‘good doctor’? a critical discourse analysis of perspectives from medical students with lived experience as patients. Medical Humanities 49 (4), pp. 613–622. Cited by: §2, §4.
  • R. L. Street Jr, G. Makoul, N. K. Arora, and R. M. Epstein (2009) How does communication heal? Pathways linking clinician–patient communication to health outcomes. Patient education and counseling 74 (3), pp. 295–301. Cited by: §2.
  • [20] A. Team Anthropic economic index report: economic primitives. Accessed: 18 Feb. 2026. External Links: Link Cited by: §2.
  • D. Tufiș and V. Barbu Mititelu (2014) The lexical ontology for romanian. In Language Production, Cognition, and the Lexicon, N. B. Nuria Gala (Ed.), Text, Speech and Language Technology, Vol. 48, pp. 491–504. Cited by: §1.
  • H. Zhou, F. Liu, B. Gu, X. Zou, J. Huang, J. Wu, Y. Li, S. S. Chen, P. Zhou, J. Liu, et al. (2023) A survey of large language models in medicine: progress, application, and challenge. arXiv preprint arXiv:2311.05112. Cited by: §2.

### v3

@@
-What Makes a Good Doctor Response? An Analysis on a Romanian Telemedicine Platform
+What Makes a Good Doctor Response? A Study on Text-Based Telemedicine
 Abstract
-Text-based telemedicine has become a common mode of care, requiring clinicians to deliver medical advice clearly and effectively in writing.
+Text-based telemedicine has become an increasingly used mode of care, requiring clinicians to deliver medical advice clearly and effectively in writing.
 As platforms increasingly rely on patient ratings and feedback, clinicians face growing pressure to maintain satisfaction scores, even though these evaluations often reflect communication quality more than clinical accuracy.
 We analyse patient satisfaction signals in Romanian text-based telemedicine.
-Using a sample of 77,334 anonymised patient question–doctor response pairs, we model feedback as a binary outcome, treating thumbs-up responses as positive and grouping negative or absent feedback into the other class.
-We extract interpretable, predominantly language-agnostic features (e.g., length, structural characteristics, readability proxies), along with Romanian LIWC psycholinguistic features and politeness/hedging markers where available.
-We train a classifier with a time-based split and perform SHAP-based analyses, which indicate that patient and clinician history features dominate prediction, functioning as strong priors, while characteristics of the response text provide a smaller but, crucially, actionable signal.
-In subgroup correlation analyses, politeness and hedging are consistently positively associated with patient feedback, whereas lexical diversity shows a negative association.
-Keywords: telemedicine; patient feedback; doctor–patient communication; Romanian NLP; interpretability; SHAP; psycholinguistic features;
+Using a sample of anonymised text-based telemedicine consultations, we model feedback as a binary outcome, treating thumbs-up responses as positive and grouping negative or absent feedback into the other class.
+We extract from doctor responses interpretable, predominantly language-agnostic features (e.g., length, structural characteristics, readability proxies), along with Romanian LIWC psycholinguistic features and politeness/hedging markers where available.
+We train a classifier with a time-based split and perform SHAP-based analyses, which indicate that metadata dominates prediction, functioning as a strong prior, while characteristics of the response text provide a smaller but actionable signal.
+In subgroup correlation analyses, politeness and hedging are consistently associated with positive patient feedback, whereas lexical diversity shows a negative association.
+Keywords: telemedicine; patient feedback; doctor–patient communication; Romanian NLP; interpretability; SHAP; psycholinguistic features.
 […]
-What Makes a Good Doctor Response?
-An Analysis on a Romanian Telemedicine Platform
+What Makes a Good Doctor Response? A Study on Text-Based Telemedicine
 Adrian Cosma, Cosmin Dumitrache, Emilian Radoi
@@
 1.  Introduction
-Text-based telemedicine has quickly evolved from a niche offering to a common way people receive care.
+Text-based telemedicine has gradually evolved from a niche offering to a more widely used way of receiving care.
 As more consultations move online, clinicians are increasingly practising medicine in writing.
@@
 In contrast to in-person care, where feedback is relatively rare and often informal, online platforms can generate large volumes of explicit evaluations for individual clinicians.
-While review systems have become ubiquitous across online platforms, applying them to medical consultations presents unique challenges.
+While review systems have become ubiquitous across online platforms, applying them to telemedicine consultations presents unique challenges.
 Doctors find themselves evaluated not by medical peers, but by patients who generally lack the clinical expertise to assess medical accuracy.
@@
 (2024).
-Most NLP research in healthcare has focused on English and other high-resource languages, with relatively limited work on Romanian.
+Most Natural Language Processing (NLP) research in healthcare has focused on English and other high-resource languages, with relatively limited work on Romanian.
 As a result, key NLP resources for Romanian, such as clinical-domain models and standardised preprocessing pipelines, remain less developed and less thoroughly validated Tufiș and Barbu Mititelu (2014); Dumitrescu et al.
 (2018).
-To keep the analysis interpretable and realistic, we therefore rely primarily on language-agnostic, surface-level measurements (e.g., length and structure) and incorporate Romanian-specific resources through the Romanian LIWC lexicon Dudău and Sava (2022).
-In this work, we use fully anonymised real-world data consisting of Romanian-language text-based medical consultations, to study which factors correlate with patient satisfaction, as reflected in post-consultation ratings.
+To keep the analysis interpretable and realistic, we therefore rely primarily on language-agnostic, surface-level measurements (e.g., length and structure) and incorporate Romanian-specific resources through the Romanian Linguistic Inquiry and Word Count (LIWC) lexicon Dudău and Sava (2022); Pennebaker et al.
+(2001).
+In this work, we use completely anonymised data consisting of Romanian-language text-based telemedicine consultations, to study which factors correlate with patient satisfaction, as reflected in post-consultation ratings.
+The data underwent irreversible de-identification to ensure that individuals could not be re-identified.
+Instead of retaining personally identifiable information, individuals were categorized into broad, non-identifiable groups based on general characteristics and prior activity.
+These groupings include distinctions such as new versus returning users, whether they have previously provided positive feedback or no feedback, and gender.
+By relying solely on aggregated and non-specific attributes, this approach preserves data utility while maintaining strict privacy protection.
 We treat feedback as the result of multiple interacting components:
   1. (i)
-    Patient profile and prior behaviour on the platform
+    Patient group information
   2. (ii)
-    Clinician profile and platform activity
+    Clinician public profile and platform activity
   3. (iii)
-    Question metadata (e.g., time, medical specialization, inter alia)
+    Question metadata (e.g., daytime interval, day of week, chosen medical specialty)
   4. (iv)
-    Properties of the written response
-Importantly, we focus on observable, surface-level attributes of doctors’ responses, features that can be measured and acted upon, rather than attempting to assess medical correctness.
-We extract features related to readability and form (e.g., length and structure), psycholinguistic signals (e.g., LIWC categories), and metadata linked to the patient, doctor and question.
+    Linguistic properties of the doctor response
+Importantly, we do not attempt to assess medical correctness, we strictly focus on observable, surface-level attributes of doctors’ responses, features that can be measured and acted upon.
+We also do not include the patient question text in our analysis.
+We extract features related to readability and form (e.g., length and structure), psycholinguistic signals (e.g., LIWC categories), and metadata linked to the patient groups, doctors and questions.
 We then train a gradient-boosting model to predict patient feedback, choosing it as a strong and simple baseline that supports post-hoc interpretability.
-To understand what the model learns, we use SHAP Lundberg and Lee (2017) to quantify global feature importance and provide local explanations for individual interactions, highlighting which response characteristics most strongly drive predicted satisfaction and how these effects vary across patient and doctor subgroups.
+To understand what the model learns, we use SHAP Lundberg and Lee (2017) to quantify global feature importance and provide local explanations for individual interactions, highlighting which response characteristics most strongly influence predicted satisfaction and how these effects vary across patient and doctor subgroups.
 2.  Related Work
@@
 (2024).
-Reliability can also be user-dependent: model behavior may degrade in public-facing use, plausibly because outputs adapt to the quality and specificity of user inputs Team ; Bean et al.
+Reliability can also be user-dependent: model behaviour may degrade in public-facing use, plausibly because outputs adapt to the quality and specificity of user inputs Team ; Bean et al.
 (2025).
-Beyond LLMs, medical humanities and sociolinguistics emphasize that patients value empathy, clarity, attention to lived experience, and appropriate boundary-setting Stergiopoulos and Martimianakis (2023).
+Beyond LLMs, medical humanities and sociolinguistics emphasize that patients value empathy, clarity, attention to lived experience and appropriate boundary-setting Stergiopoulos and Martimianakis (2023).
 A well presented medical advice is more likely to have a positive impact Martin et al.
@@
 Rather than asking whether LLMs provide correct answers, we study what patients reward in human clinician writing.
-Using a large Romanian telemedicine dataset (77k+ question–answer pairs), we measure which observable properties of replies (surface form, readability proxies, psycholinguistic features) are associated with positive feedback.
-This complements Romanian-focused efforts on optimizing LLM-assisted doctor communication that rely on smaller exploratory datasets Niculae et al.
+We measure which observable properties of doctor replies (surface form, readability proxies, psycholinguistic features) are associated with positive feedback.
+This complements Romanian-focused efforts on optimizing LLM-assisted doctor communication Niculae et al.
 (2025a).
@@
 3.  Method
-We study text-based consultations consisting of patient questions paired with doctor responses (N=77,334N{=}77{,}334).
-After an interaction, patients may leave structured feedback (thumbs-up / thumbs-down), but in practice most interactions receive no explicit rating and explicit negative feedback is rare.
+We study anonymised text-based consultations.
+After an interaction, which consists of a patient question and a doctor response, patients may leave structured feedback (thumbs-up / thumbs-down).
+In practice most interactions receive no explicit rating and explicit negative feedback is rare.
 We therefore model patient satisfaction as a binary prediction task, treating thumbs-up as the positive class and grouping thumbs-down and missing feedback into the other class.
-The resulting dataset is approximately balanced (39,201 positive vs.
-38,133 other).
-Although fully anonymised, the data is private and cannot be released.
+As such, the task is to predict whether a doctor response will receive positive feedback or not.
+This formulation roughly corresponds to the Positive-Unlabeled learning (PU-learning) Bekker and Davis (2020) setup.
+The patient and doctor usually do not have a prior connection and each interaction is stand-alone.
+Patients can either choose which doctor to send their question to, or they can choose to have the doctor assigned by the platform.
+Although completely anonymised, the data is privately owned and cannot be released.
 A particular aspect of the studied text consultations is that doctors may address clarification questions before providing an answer.
@@
 Feature extraction To keep the analysis interpretable and robust in a lower-resource language setting, we focus primarily on surface-level, mostly language-agnostic features, and add Romanian-specific resources where available.
-We extract features from four sources: (i) patient profile – platform tenure, activity, and prior behaviour (e.g., historical thumbs-up rate, number of posed questions per year), (ii) doctor profile – metadata (e.g., workplace, education) and platform activity and historical feedback rate.
-(iii) question metadata and patient text – basic structure and readability proxies (e.g., length), plus LIWC psycholinguistic categories, and (iv) doctor response text – length / structure (characters, tokens, punctuation counts such as question marks), concreteness proxies (e.g., number of digits), and lexicon-based markers (e.g., politeness, hedging, recommendation cues), as well as LIWC features.
-We compute a fixed feature vector of 250 features spanning metadata, history and text-derived signals.
-Prediction model and evaluation protocol We train a gradient-boosted decision tree classifier for binary feedback prediction.
+We extract features from four sources: (i) patient group information – new vs returning user, prior activity (e.g., previous thumbs-up), (ii) public doctor profile – metadata (e.g., workplace, education), platform activity and historical feedback rate, (iii) question metadata (e.g., daytime interval, day of week, chosen medical specialty), (iv) doctor response text – binned quantities analysed in ranges for length / structure (number of characters, tokens, punctuation such as question marks), concreteness proxies (e.g., number of digits), and lexicon-based markers (e.g., politeness, hedging, recommendation cues), as well as LIWC features.
+We compute a fixed feature vector, which is used for predictive modelling.
+Prediction model and evaluation protocol We train a gradient-boosted decision tree classifier (CatBoost) for binary feedback prediction.
 To reduce temporal leakage, we use a time-based split: training uses earlier interactions and evaluation is performed on later interactions, for a 80:20 train:test split.
@@
 We summarize global importance with mean absolute SHAP values and use local explanations to inspect individual decisions.
-In addition, to probe whether associations differ across populations, we perform subgroup correlation analyses between selected text features and feedback, stratifying by patient gender (patient preferences) and doctor gender (doctor behavior), and report the resulting patterns descriptively.
+In addition, to probe whether associations differ across populations, we perform subgroup correlation analyses between selected text features and feedback, stratifying by patient gender (patient preferences) and doctor gender (doctor behaviour), and report the resulting patterns descriptively.
 4.  Results
@@
 New Patients Val 0.6479 0.6003 0.5804 0.6061 0.5963
-Old Patients Val 0.7656 0.6533 0.6951 0.6968 0.6937
+Returning Patients Val 0.7656 0.6533 0.6951 0.6968 0.6937
 Simple Questions Val High-conf 0.6951 0.6138 0.6456 0.7724 0.6416
@@
 All Questions Val High-conf 0.8357 0.6987 0.7889 0.8001 0.7806
-Table 1: Feedback prediction performance across question subsets and patient tenure groups (New vs. Old), reported on the validation split and a high-confidence subset.
-Figure 1: Mean absolute SHAP values for the CatBoost feedback classifier trained on 250 features spanning patient profile/history, doctor profile/history and text-derived features from both sides.
+Table 1: Feedback prediction performance across question subsets and patient groups (New vs. Returning), reported on the validation split and a high-confidence subset.
+Figure 1: Mean absolute SHAP values for the CatBoost feedback classifier.
 Larger values indicate greater impact on predictions (magnitude only; SHAP sign is not shown).
+Quantities are binned and are not processed in absolute numbers.
 Figure 2: Pearson correlations between selected doctor-response text features and positive feedback, aggregated over all interactions.
-Positive values indicate features that tend to co-occur with thumbs-up feedback, while negative values indicate features that tend to co-occur with negative or missing feedback.
+Positive values indicate features that tend to co-occur with thumbs-up feedback, while negative values indicate features that tend to co-occur with thumbs-down or missing feedback.
+Quantities are binned and are not processed in absolute numbers.
 In Table 1, we show classifier performance across subgroups.
@@
 Correlation analysis by doctor and patient gender.
-Figure 2 shows Pearson correlations between doctor-text features and binary positive feedback, computed separately by doctor gender (Doctor Behaviour) and by patient gender (Patient Preferences).
-Across all splits, the largest positive correlations are for politeness markers and hedging, consistent with prior work Stergiopoulos and Martimianakis (2023).
-Smaller positive correlations appear for features that often signal concrete, actionable replies, such as the number of digits, explicit recommendations, and question marks.
+Figure 2 shows Pearson correlations between doctor-text features and binary feedback, computed separately by doctor gender (Doctor Behaviour) and by patient gender (Patient Preferences).
+Across all splits, the largest positive correlations are for politeness markers and hedging in the doctor response text, consistent with prior work Stergiopoulos and Martimianakis (2023).
+Smaller positive correlations appear for features that often signal concrete, actionable replies, such as the number of digits, explicit recommendations and question marks.
 The most consistent negative correlation is for lexical diversity, with additional negative correlations for several LIWC categories (e.g., liwc_work, liwc_affiliation, liwc_social).
 Higher lexical diversity can reflect more complex, less standardized phrasing, which may increase cognitive load; conversely, simpler, more formulaic language may read as clearer and safer Stergiopoulos and Martimianakis (2023).
-The overall pattern is similar across groups, suggesting that both male and female patients tend to respond positively to polite, cautious, and concrete messages.
+The overall pattern is similar across the two gender groups, suggesting that both male and female patients tend to respond positively to polite, cautious and concrete messages.
 Differences by patient gender are small.
-Differences by doctor gender are more visible: for several of the top positive features (especially politeness, hedging and digits), correlations are larger within the subset of replies written by female doctors than within those written by male doctors.
+Differences by doctor gender are more visible: for several of the top positive features (especially politeness, hedging and number of digits), correlations are larger within the subset of replies written by female doctors than within those written by male doctors.
 Correlation magnitudes are modest overall (roughly $|r|\approx 0.02$$0.15$), so these results should be treated as descriptive.
@@
 Figure 1 reports mean absolute SHAP values (feature importance).
-Overall, the model relies most on patient history features.
-The largest contributions come from the previous thumbs-up rate and the previous number of questions.
-These features mainly act as strong priors: patients with a higher past tendency to give positive feedback (or higher platform activity/tenure) are easier to predict.
-Doctor-level features contribute less than patient-level history.
-The top doctor predictors are previous received thumbs-up rate, doctor workplace, number of previously helped patients, and number of years on the platform, along with past positive / negative feedback rates.
-This suggests the model captures persistent differences between clinicians and settings (e.g., workplace) beyond what is in the text.
-Text features have smaller mean $|\mathrm{SHAP}|$ values than profile/history features, indicating that metadata dominates the prediction.
-Within text, doctor-side features are more influential than patient-side features.
+Overall, the model relies most on patient group information features.
+The largest contributions come from the returning users feature and the previous thumbs-up feature.
+These features mainly act as strong priors: patients that previously gave positive feedback are more predictable.
+Doctor-level features contribute less than patient group information.
+The top doctor predictors are previously received thumbs-up rate, doctor workplace, number of previously helped patients, and number of years on the platform, along with past positive / negative feedback rates.
+This suggests the model captures persistent differences between clinicians and settings (e.g., workplace) beyond the doctor response.
+Doctor response text features have smaller mean $|\mathrm{SHAP}|$ values than patient groups and doctor public profiles features, indicating that metadata dominates the prediction.
 The most important doctor-text signals include number of questions, number of digits present, and several LIWC categories (e.g., liwc_work, liwc_negemo, liwc_time), as well as length proxies (n_characters, n_tokens).
@@
 Digits often encode dosage, timing, thresholds or step-by-step plans, which are surface markers of actionability.
-Patient-text features have comparatively small contributions.
-Satisfaction signals are identity- and expectation-loaded: many ratings reflect the rater’s baseline tendency and the clinician’s reputation/context more than the linguistic content.
+Satisfaction signals are identity- and expectation-loaded, as ratings reflect the rater’s baseline tendency and the clinician’s reputation/context more than the response.
 This is consistent with "perception-first" evaluation dynamics discussed in prior work Shekar et al.
 (2024), and motivates controlling for priors before making communication recommendations.
+Figure 3: Mean absolute SHAP values for the most important doctor text features, computed separately for each medical question specialty.
+Bars show the response-text features with the largest contribution magnitude to the feedback prediction model within each specialty.
+Values indicate feature importance magnitude only and do not show effect direction.
+Quantities are binned and are not processed in absolute numbers.
+Figure 3 shows that, when restricting the analysis to doctor-response text and stratifying by medical specialty, the most predictive cues become more specialty-specific.
+Some features recur across categories, especially digits, time-related language, and task-oriented terms, suggesting that patient feedback favors responses that are concrete, procedural and actionable.
+At the same time, several categories display clinically intuitive lexical patterns: dermatology emphasizes perceptual and symptom-description language; endocrinology is strongly driven by numeric content; gastroenterology highlights ingestion- and risk-related terms; ENT uniquely surfaces hearing-related language; pediatrics gives more weight to social and second-person terms; and gynecology/urology place greater emphasis on sexual and politeness-related language.
+Overall, results suggests that positive feedback is associated not only with general clarity, but with responses whose language is appropriately matched to the communicative demands of the medical specialty.
 5.  Conclusions
-We analysed 77k anonymised Romanian patient question–doctor response pairs to examine which characteristics of written clinician replies are associated with positive patient feedback in text-based telemedicine consultations.
-The main finding is that satisfaction is driven largely by stable priors (patient and doctor history), so predicting feedback mostly captures who tends to rate positively and who tends to receive positive ratings.
-Still, the response text contains smaller but actionable signals: politeness and calibrated uncertainty are consistently associated with positive feedback, as are concrete, action-oriented cues (e.g., digits and recommendations), while higher lexical diversity correlates negatively.
-Presumably, more high level and semantic response features metrics tailored to this specific use-case can be discovered automatically using feature discovery methods Cosma et al.
+We analysed anonymised Romanian text-based telemedicine consultations to examine which characteristics of written clinician replies are associated with positive patient feedback.
+The main finding is that satisfaction is driven largely by stable priors (patient group information and doctor activity), so predicting feedback mostly captures who tends to rate positively and who tends to receive positive ratings.
+Still, the response text contains smaller but actionable signals: politeness and hedging are consistently associated with positive feedback, as are concrete, action-oriented cues (e.g., digits and recommendations), while higher lexical diversity correlates negatively.
+Presumably, more high-level and semantic response features tailored to this specific use-case could be discovered automatically using feature discovery methods Cosma et al.
 (2026).
-These results suggest that patient feedback should be interpreted primarily as an incentive for clear, empathetic written communication, rather than as a proxy for medical accuracy.
+The results suggest that patient feedback should be interpreted primarily as a signal of clear, empathetic written communication, rather than as a proxy for medical accuracy.
+Future work should focus on modelling higher-level, semantic features of the doctor responses, such as response clarity and whether the response contains adequate explanations where appropriate.
+These aspects are not easily captured through lexicon-based approaches and may require the use of pretrained language models Cosma et al.
+(2026).
+6.  Bibliographical References
   • A.
@@
 Cited by: §2.
-  • A. Cosma, O. Szehr, D. Kletz, A. Antonucci, and O. Pelletier (2026) Automatic prompt optimization for dataset-level feature discovery. External Links: 2601.13922, Link Cited by: §5.
+  • J. Bekker and J. Davis (2020) Learning from positive and unlabeled data: a survey. Machine learning 109 (4), pp. 719–760. Cited by: §3.
+  • A. Cosma, O. Szehr, D. Kletz, A. Antonucci, and O. Pelletier (2026) Automatic Prompt Optimization for Dataset-Level Feature Discovery. External Links: Link, 2601.13922 Cited by: §5, §5.
   • D.
@@
 1780–1792.
-External Links: Link, Document, ISBN 979-8-89176-333-3 Cited by: §2.
-  • A. Niculae, A. Cosma, and E. Radoi (2025b) A retrieval-based approach to medical procedure matching in romanian. arXiv preprint arXiv:2503.20556. Cited by: §1.
+External Links: Document, ISBN 979-8-89176-333-3, Link Cited by: §2.
+  • A.
+Niculae, A.
+Cosma, and E.
+Radoi (2025b) A Retrieval-Based Approach to Medical Procedure Matching in Romanian.
+In Proceedings of the 24th Workshop on Biomedical Language Processing, D.
+Demner-Fushman, S.
+Ananiadou, M.
+Miwa, and J.
+Tsujii (Eds.), Viena, Austria, pp.
+167–175.
+External Links: Document, ISBN 979-8-89176-275-6, Link Cited by: §1.
   • H.
@@
 External Links: Document, Link Cited by: §1.
+  • J. W. Pennebaker, M. E. Francis, R. J. Booth, et al. (2001) Linguistic inquiry and word count: liwc 2001. Mahway: Lawrence Erlbaum Associates 71 (2001), pp. 2001. Cited by: §1.
   • M. Reis, F. Reis, and W. Kunde (2024) Influence of believed AI involvement on the perception of digital medical advice. Nature Medicine 30 (11), pp. 3098–3100. Cited by: §2.
@@
 I.
-Ionescu (2025) MedQARo: a large-scale benchmark for evaluating large language models on medical question answering in romanian.
-arXiv preprint arXiv:2508.16390.
-Cited by: §1.
+Ionescu (2026) A large-scale benchmark for evaluating large language models on medical question answering in Romanian.
+npj Digital Medicine.
+External Links: Document, ISBN 2398-6352, Link Cited by: §1.
   • J.
@@
 Barbu Mititelu (2014) The Lexical Ontology for Romanian.
-In Language Production, Cognition, and the Lexicon, N.
-B.
-Nuria Gala (Ed.), Text, Speech and Language Technology, Vol.
+In Language Production, Cognition, and the Lexicon, Text, Speech and Language Technology, Vol.
 48, pp.